OpenStudy (anonymous):

please help.! Two legs of a triangle measure 16 m and 34 m. How many meters long is the hypotenuse? Round your answer to the nearest tenth.

OpenStudy (anonymous):

familiar with the Pythagorean theorem?

hero (hero):

16^2 + 34^2 = h^2

OpenStudy (anonymous):

let the hypoth. be C, A=16, and B=34 \[A^2+B^2=C^2\] and therefore: \[C= \sqrt{A^2+B^2}\] C= \[\sqrt{256+1156}\] C=37.57658

OpenStudy (anonymous):

make sure you learn how to do the question. bye